Richetti produces stereo speakers. The selling price per pair of speakers is $1,800. There is no beginning inventory. Costs involved in production are: Direct material $161 Direct labor 183 Variable manufacturing overhead 107 Total variable manufacturing costs per unit $451 Fixed manufacturing overhead per year $594,000 In addition, the company has fixed selling and administrative costs: Fixed selling costs per year $200,500 Fixed administrative costs per year $102,400 During the year, Richetti produces 1,500 pairs of speakers and sells 1,200 pairs.

What is cost of goods sold using full costing?

Cost of goods sold $__________
Answer #1

Answer : $ 1,016,400

Units Cost Per Unit Under Fullcosting

= Variable Manufacturing Cost per Unit + Fixed Manufacturing cost Per Unit

= $451 + ( 594000/1500) = 451 +396= $847

Cost of Goods Sold = Units Sold * Prodcut Cost Per Unit

Cost of Goods Sold = 1200*847= $ 1,016,400 (Answer)